Locate the centroid (7, ý) of this area using BOTH (a) integration and (b) the Composite Method. The x-coordinate d is 5 m.
Answer #1

y=- 2.5x +20 20mA ec t - DE dz 7.5 di (5,0) at point A y=pom ,x=0 at point B 20 =5m, y = -2.5(5)+20 = -12.5+ 20 = BE (5,7.5)

considering 5 = xy da Suda x(-2.5x+20)dx X = 842.5x+20) de I €2.5x2 +20x) dx (-2.50 +20) dx - 2.5x3 20x295 Z => 1 [-2.523+27- JE [ 25*8 logettoox7 [ 2552 + 20s]: = ý = I (6.25 (53_10o (5²) + 400s)) [ - 2.5(52+2015] - ū= [960.4166–1950 + povo] ? [-3150 = y = AYIt AzY2 A, 3G + A₂X2 Ait Az At Az - - - 31.254 1.66) +(37.5725) = 2.12 m. 31.25 + 37.5 D } = (31.25x11.66) +(37.5%